OpenStudy (anonymous):

derivative of 2x-3)^4 (x^2+x+1)^5

OpenStudy (anonymous):

let y=(2x-3)^4(x^2+x+1)^5 either use f'(x)=uv'+u'v or take the log of both sides and diff.

OpenStudy (anonymous):

i dont understand. can you please show me how

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=4(2x-3)^3*2*(x^2+x+1)^5+(2x-3)^4*5(x^2+x+1)^4(2x+1)\] simplify

OpenStudy (anonymous):

or log y=4log(2x-3)+5 log(x^2+x+1) diff.